OpenStudy (anonymous):

The table shows the amount of red paint that can be added to different amounts of white paint while making a custom paint color: Paint Color Quarts of Red Paint Quarts of White Paint 4 20 5 25 7 35 8 40 Which statement is true about the ratio of red paint to white paint in the paint color? It is 1:16. It is 16:1. It is 5:1. It is 1:5.

OpenStudy (anonymous):

@dimensionx

OpenStudy (anonymous):

@dimensionx are you there?

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

Write a proportion to represent the values and solve for the ratio in the table. Red Paint 4, White Paint 20. \[\frac{ 4 }{ 20 } = ?\]

OpenStudy (anonymous):

what would be the next step?

OpenStudy (anonymous):

Divide, 4/20

OpenStudy (anonymous):

or 5/25 or 7/35 or 8/40

OpenStudy (anonymous):

= 5

OpenStudy (anonymous):

Try again

OpenStudy (anonymous):

4 divided by 20?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

.2

OpenStudy (anonymous):

Think of it this way: \[\frac{ 4\div4 }{ 20 \div 4 } = \frac{ ? }{ ? }\]

OpenStudy (anonymous):

1.2

OpenStudy (anonymous):

Numerator: What is 4 divided by 4? Denominator: What is 20 divided by 4?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

4 divided by 4 is 1, correct. But try again with 20 divided by 4.

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

yes, so you have a 1:5 ratio of Red:White paint. For every 4 cans of red paint, you need 20 cans of white paint. SO you can simplify: For every 1 can of red paint, you'll need 5 cans of white paint.

OpenStudy (anonymous):

oh so then the ratio is 1:5. choice d

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks so much! can you check if this is correct?

OpenStudy (anonymous):

No, it is not a proportional relationship because the graph is not in a straight line

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok, thanks! i gave you a medal